Project

General

Profile

NDOS Commissioning Meeting 2011/09/22

Attendance: Jon, Martin, Mat, Zukai, Gavin, Jarek, Jan, Enrique, Rafael, Susan, Zeb, Nick, Minerba,

Agenda:

  1. Beam & Detector Update
  2. DAQ Status & Plans
  3. DCS Status & Plans
  4. AOB

Beam & Detector Update

  • NuMI target NT07 has been installed and they are currently doing target scans.
  • Next Tuesday (9/27/11), Walter Jaskierny will be removing his power monitoring equipment from rack 2 in the NDSB. This will require powering down that rack, thus the HV system. Jon will try to arrange for this to happen at 9 am. Martin: Gennadiy should be available in case there are problems. Jon will follow up with Peter
  • Mat: Martin and Mat went out to detector after commissioning meeting to investigate 3 FEBs (none of which had APDs) that were faulty. Two were not responding to any commands, the other was responding in DSO scans but was reporting FEB temp. alarm (temp=0). FEBs were located on DCMs 2-2, 2-3, 4-2. One had blown fuse which Martin replaced. One FEB was corroded due to water. Nothing was obviously wrong with the other FEB. The two FEBs that were removed were sent to Harvard for further investigation.
  • Mat: also adjust swinging door sensor so sporadic alarms should go away.

DCS Update

  • Martin:
    • current monitoring on DCS power supply monitor is now displayed in the Labview program.
    • if we start cooling next week, contact Gennadiy who knows how to re-enable cooling monitoring

DAQ Update

  • None

AOB

  • Jarek is working on a Nearline Checklist for shifters. A preliminary version will likely be available for shifters to fill out this weekend.